OpenStudy (anonymous):

9) Find three consecutive odd numbers such that the square of the second number is 192 less than the square of the third number.

hero (hero):

Three consecutive odd numbers means 1st Number = x 2nd Number = x + 2 3rd Number = x + 4 Now put those into an equation under the given restrictions: \((x+2)^2 = (x+4)^2 - 192\)

OpenStudy (anonymous):

Thank You, I have a really hard time with word problems

hero (hero):

yw

hero (hero):

Did you find the three numbers yet?

OpenStudy (anonymous):

I think they are 49, 47, and 45